# RoSTE: An Efficient Quantization-Aware Supervised Fine-Tuning Approach for Large Language Models

### v1

@@ Abstract
Supervised fine-tuning is a standard method for adapting pre-trained large language models (LLMs) to downstream tasks. Quantization has been recently studied as a post-training technique for efficient LLM deployment. To obtain quantized fine-tuned LLMs, conventional pipelines would first fine-tune the pre-trained models, followed by post-training quantization. This often yields suboptimal performance as it fails to leverage the synergy between fine-tuning and quantization. To effectively realize low-bit quantization of weights, activations, and KV caches in LLMs, we propose an algorithm named Rotated Straight-Through-Estimator (RoSTE), which combines quantization-aware supervised fine-tuning (QA-SFT) with an adaptive rotation strategy that identifies an effective rotation configuration to reduce activation outliers. We provide theoretical insights on RoSTE by analyzing its prediction error when applied to an overparameterized least square quantized training problem. Our findings reveal that the prediction error is directly proportional to the quantization error of the converged weights, which can be effectively managed through an optimized rotation configuration. Experiments on Pythia and Llama models of different sizes demonstrate the effectiveness of RoSTE. Compared to existing post-SFT quantization baselines, our method consistently achieves superior performances across various tasks and different LLM architectures.
@@ 1 Introduction
We conduct experiments on fine-tuning Pythia and Llama models, demonstrating the effectiveness of RoSTE. Overall, we believe this is the first work that develops an efficient quantization algorithm for the SFT process, with theoretical justification and practical effectiveness.
@@ 2.2 Quantization-Aware Training (QAT)
for any differentiable function $g$. This approximation utilizes the insight that a quantizer behaves like an identity function in low resolution, while tolerating a gradient bias since quantization error persists in high resolution. In practice, STE is known to work well in training quantized neural network models (Li et al., 2017; Yin et al., 2019) as well as LLMs (Liu et al., 2023).
@@ 4 Theoretical Insights of RoSTE
Suppose that the QAT stage of RoSTE is run with $T\gg 1$ inner-loop iterations. Applying the theorem shows that given $\overline{\mathbf{R}}$, the resultant prediction error of the intermediate model $\mathbf{w}^{T}$ will be bounded by ${\cal O}(\sum_{s=0}^{T}\left(1-\mu\right)^{T-s}\mathbb{E}\left[{\|Q_{w}(\overline{\mathbf{R}}\mathbf{w}^{s})-\overline{\mathbf{R}}\mathbf{w}^{s}\|_{\mathbf{G}}^{2}}\right])$, i.e., a weighted sum of the weight quantization errors during the QAT process. Due to the exponential term $\left(1-\mu\right)^{T-s}$, the prediction error is dominated by the weight quantization error of recent iterates. Crucially, the above analysis shows that the rotation matrices play a pivoting role in the performance of QAT. This inspires us to apply ${\cal E}(\cdot)$ in (12) to guide us in the selection for optimal rotation matrices.
@@
Table 1: Results on Exp.1. Accuracies of the 4-bit quantized Pythia 6.9B model fine-tuned using the Reddit TL;DR dataset. FP16 refers to using 16 bits half-precision floating points, and W4A4KV4 refers to using 4 bit quantizations on weights, activation, and KV cache.
[…]
Table 2: Results on Exp.2. Accuracies of the 4-bit quantized Llama 3.1 8B model fine-tuned on the Tulu 3 SFT mixture dataset. BP16 refers to using 16-bit brain floating points format, and W4A4KV4 refers to using 4-bit quantizations on weights, activation, and KV cache.
[…]
@@ 5 Experiments
We evaluate the performance of the proposed RoSTE algorithm for QA-SFT on two standard sets of open-source models and datasets. For the first experiment (Exp.1), we fine-tune the pre-trained Pythia 1B/6.9B models (Biderman et al., 2023) on the Reddit TL;DR Summarization dataset (Huang et al., 2024) with evaluation on the TL;DR test dataset using the ROUGE metric (Lin, 2004). For the second experiment (Exp.2), we fine-tune the pre-trained Llama 3.1 8B model (Dubey et al., 2024) on the Tulu 3 SFT mixture dataset (Lambert et al., 2024) with real-world downstream task evaluations (Gao et al., 2021). These tasks include TruthfulQA (Lin et al., 2021), MMLU-Pro (Wang et al., 2024b), BigBenchHard (Suzgun et al., 2022), AGIEval (Zhong et al., 2023), GSM8K (Cobbe et al., 2021), and MATH (Hendrycks et al., 2020).
For the RoSTE algorithm, while we relaxed the lower level as a $\ell$-variable binary combinatorial problem (9), solving this sub-problem has a complexity of $\mathcal{O}(2^{\ell})$ which is still intractable for models like Llama 3.1 8B with $\ell=3\times 32+1$. As a remedy, we estimate the solution of (9) by sharing the rotation matrices across different layers. This reduces the problem into a 4-variable binary combinatorial optimization. Our experiment results suggest that the above shared variable approximation suffices to find rotation matrices that effectively reduce outliers. Lastly, we set $K=1$ where a one-shot rotation configuration adaptation by pre-trained model is found to perform well. We anticipate the performance to further improve with increasing $K$. More implementation details can be found in Appendices C, D.
@@ 5.1 Accuracy of Quantized Models
Lastly, in the appendix, we provide additional results for Pythia 1B and W4A8KV4 in Table 7, and for Llama 3.1 8B with W4A8KV4 quantization in Table 8.
@@ 5.2 Ablation Study
Secondly, we take a closer look at the effects of outlier reduction to justify our claims on the use of random Walsh-Hadamard rotation matrices. Fig. 2 compares the distribution of the input activations of fine-tuned models trained by STE and RoSTE at layer 30. We observe that the model produced by RoSTE exhibits no activation outliers, while STE suffers from activation outliers even at convergence. For a more detailed and comprehensive comparison, see Fig. 5 and 6 in the appendix. Furthermore, Fig. 3 shows the trajectory of the quantization error (12) during training. As expected, we see that the quantization error of RoSTE is much lower than that in STE, thus suggesting a lower bias in the solution obtained [cf. Theorem 4.3].
@@ Appendix C Details of Experiment Settings
Model Pythia 1B Pythia 6.9B Llama 3.1 8B
Epoch 1 1 2
Batch Size (Per GPU) 16 1 1
Gradient Accumulation 1 16 16
Optimizer AdamW AdamW AdamW
Learning Rate 3e-5 3e-5 5e-6
LR Schedule cosine cosine linear
Warmup Ratio 0 0 0.03
Max. Seq. Length 2048 2048 1024
# Training Samples 117k 117k 100k
@@
Method QA-SFT (i.e., STE or RoSTE)
Model Pythia 1B Pythia 6.9B Llama 3.1 8B
Epoch 1 1 2
Batch Size (Per GPU) 16 1 1
Gradient Accumulation 1 16 16
Optimizer AdamW AdamW AdamW
Learning Rate {3e-5, 6e-6, 3e-6} {3e-5, 6e-6, 3e-6} {5e-6, 1e-6, 5e-7}
LR Schedule cosine cosine linear
Warmup Ratio 0 0 0.03
Max. Seq. Length 2048 2048 1024
# Training Samples 117k 117k 100k
clipping factor {1, 0.95, 0.9} {1, 0.95, 0.9} {1, 0.95, 0.9}
@@ Appendix F Additional Experiments
Table 7: Additional experiments for Exp.1 with different bit-width configurations and different model sizes. W4A8KV4 refers to using 4-bit quantizations on weights, 8-bit on activation, and 4-bit on KV cache.
Method ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSum ROUGE (Avg.)
@@
Table 8: Additional experiments for Exp.2 on different bit-width configurations and different model size. W4A8KV4 refers to using 4 bit quantizations on weights, 8 bit on activation, and 4 bit on KV cache.
Method TruthfulQA MMLU-Pro BigBenchHard AGIEval GSM8K Math Avg.

### v2

@@ Abstract
Supervised fine-tuning is a standard method for adapting pre-trained large language models (LLMs) to downstream tasks. Quantization has been recently studied as a post-training technique for efficient LLM deployment. To obtain quantized fine-tuned LLMs, conventional pipelines would first fine-tune the pre-trained models, followed by post-training quantization. This often yields suboptimal performance as it fails to leverage the synergy between fine-tuning and quantization. To effectively realize low-bit quantization of weights, activations, and KV caches in LLMs, we propose an algorithm named Rotated Straight-Through-Estimator (RoSTE), which combines quantization-aware supervised fine-tuning (QA-SFT) with an adaptive rotation strategy that identifies an effective rotation configuration to reduce activation outliers. We provide theoretical insights on RoSTE by analyzing its prediction error when applied to an overparameterized least square quantized training problem. Our findings reveal that the prediction error is directly proportional to the quantization error of the converged weights, which can be effectively managed through an optimized rotation configuration. Experiments on Pythia, Qwen and Llama models of different sizes demonstrate the effectiveness of RoSTE. Compared to existing post-SFT quantization baselines, our method consistently achieves superior performances across various tasks and different LLM architectures. Our code is available at https://github.com/Hong-Lab-UMN-ECE/RoSTE.
Figure 1: RoSTE surpasses the performance of SOTA quantization methods on fine-tuning benchmark. Horizontal axis represents the total amount of hours needed to fine-tune pre-trained LLMs on a server of 8 $\times$ A100 NVIDIA GPUs.
@@ 1 Introduction
We conduct experiments on fine-tuning Pythia, Qwen and Llama models, demonstrating the effectiveness of RoSTE. An accuracy-vs-training-time plot in Figure 1 illustrates the practical benefit of RoSTE for obtaining a fine-tuned quantized LLM. Overall, we believe this is the first work that develops an efficient quantization algorithm for the SFT process, with theoretical justification and practical effectiveness.
@@ 2.2 Quantization-Aware Training (QAT)
for any differentiable function $g$. This approximation utilizes the insight that a quantizer behaves like an identity function in low resolution, while tolerating a gradient bias since quantization error persists in high resolution. In practice, STE is known to work well in training quantized neural network models (Li et al., 2017; Yin et al., 2019) as well as LLMs (Liu et al., 2023; Panferov et al., 2025).
@@ 4 Theoretical Insights of RoSTE
Suppose that the QAT stage of RoSTE is run with $T\gg 1$ inner-loop iterations. Applying the theorem shows that given $\overline{\mathbf{R}}$, the resultant prediction error of the intermediate model $\mathbf{w}^{T}$ will be bounded by ${\cal O}(\sum_{s=0}^{T}\left(1-\mu\right)^{T-s}\mathbb{E}\left[{\|Q_{w}(\overline{\mathbf{R}}\mathbf{w}^{s})-\overline{\mathbf{R}}\mathbf{w}^{s}\|_{\mathbf{G}}^{2}}\right])$, i.e., a weighted sum of the weight quantization errors during the QAT process. Due to the exponential weighting $\left(1-\mu\right)^{T-s}$, the prediction error is dominated by the weight quantization error of recent iterates. Crucially, the above analysis shows that the rotation matrices play a pivoting role in the performance of QAT. This inspires us to apply ${\cal E}(\cdot)$ in (12) to guide us in the selection for optimal rotation matrices.
@@
Table 1: Results on Exp.1. Accuracies of the 4-bit quantized Pythia 6.9B and Qwen 2.5 7B models fine-tuned using the Reddit TL;DR dataset. FP16 and BF16 refer to using 16-bit half-precision floating points and 16-bit brain floating points formats, respectively, and W4A4KV4 refers to using 4-bit quantizations on weights, activation, and KV cache.
[…]
Table 2: Results on Exp.2. Accuracies of the 4-bit quantized Llama 3.1 8B model fine-tuned on the Tulu 3 SFT mixture dataset. BF16 refers to using 16-bit brain floating points format, and W4A4KV4 refers to using 4-bit quantizations on weights, activation, and KV cache.
[…]
@@ 5 Experiments
We evaluate the performance of the proposed RoSTE algorithm for QA-SFT on two standard sets of open-source models and datasets. For the first experiment (Exp.1), we fine-tune the pre-trained Pythia 1B/6.9B models (Biderman et al., 2023) and Qwen 2.5 0.5B/7B models (Yang et al., 2024) on the Reddit TL;DR Summarization dataset (Huang et al., 2024) with evaluation on the TL;DR test dataset using the ROUGE metric (Lin, 2004). For the second experiment (Exp.2), we fine-tune the pre-trained Llama 3.1 8B model (Dubey et al., 2024) on the Tulu 3 SFT mixture dataset (Lambert et al., 2024) with real-world downstream task evaluations (Gao et al., 2021). These tasks include TruthfulQA (Lin et al., 2021), MMLU-Pro (Wang et al., 2024b), BigBenchHard (Suzgun et al., 2022), AGIEval (Zhong et al., 2023), GSM8K (Cobbe et al., 2021), and MATH (Hendrycks et al., 2020).
For the RoSTE algorithm, while we relaxed the lower level as a $\ell$-variable binary combinatorial problem (9), solving this sub-problem has a complexity of $\mathcal{O}(2^{\ell})$ which is still intractable for models like Llama 3.1 8B with $\ell=3\times 32+1$. As a remedy, we estimate the solution of (9) by computing only $\mathcal{E}({\cal W}^{kT},\{{\bf I}\}_{i=0}^{\ell-1})$ and $\mathcal{E}({\cal W}^{kT},\{{\bf H}\}_{i=0}^{\ell-1})$, then we determine each layer’s $\mathbf{R}_{i}$ by comparing the layer-wise quantization error. Lastly, we set $K=1$ where a one-shot rotation configuration adaptation by pre-trained model is found to perform well. We anticipate the performance to further improve with increasing $K$. More implementation details can be found in Appendices C, D.
@@ 5.1 Accuracy of Quantized Models
Lastly, in the appendix, we provide additional results on Pythia 1B and W4A8KV4 in Table 7, Qwen 2.5 in Table 8, and Llama 3.1 8B and W4A8KV4 quantization in Table 9.
@@ 5.2 Ablation Study
Secondly, we take a closer look at the effects of outlier reduction to justify our claims on the use of random Walsh-Hadamard rotation matrices. Fig. 3 compares the distribution of the input activations of fine-tuned models trained by STE and RoSTE at layer 30. We observe that the model produced by RoSTE exhibits no activation outliers, while STE suffers from activation outliers even at convergence. For a more detailed and comprehensive comparison, see Fig. 6 and 7 in the appendix. Furthermore, Fig. 4 shows the trajectory of the quantization error (12) during training. As expected, we see that the quantization error of RoSTE is much lower than that in STE, thus suggesting a lower bias in the solution obtained [cf. Theorem 4.3].
@@ Appendix C Details of Experiment Settings
Model Pythia 1B Pythia 6.9B Qwen2.5 0.5B Qwen2.5 7B Llama 3.1 8B
Epoch 1 1 1 1 2
Batch Size (Per GPU) 16 1 16 1 1
Gradient Accumulation 1 16 1 16 16
Optimizer AdamW AdamW AdamW AdamW AdamW
Learning Rate 3e-5 3e-5 5e-5 1e-5 5e-6
LR Schedule cosine cosine cosine cosine linear
Warmup Ratio 0 0 0 0 0.03
Max. Seq. Length 2048 2048 2048 2048 1024
# Training Samples 117k 117k 117k 117k 100k
@@
Method QA-SFT (i.e., STE or RoSTE) Model Pythia 1B Pythia 6.9B Qwen2.5 0.5B Qwen2.5 7B Llama 3.1 8B Epoch 1 1 1 1 2 Batch Size (Per GPU) 16 1 16 1 1 Gradient Accumulation 1 16 1 16 16 Optimizer AdamW AdamW AdamW AdamW AdamW Learning Rate {3e-5, 6e-6, 3e-6} {3e-5, 6e-6, 3e-6} {5e-5, 1e-5, 5e-6} {5e-5, 1e-5, 5e-6} {5e-6, 1e-6, 5e-7} LR Schedule cosine cosine cosine cosine linear Warmup Ratio 0 0 0 0 0.03 Max. Seq. Length 2048 2048 2048 2048 1024 # Training Samples 117k 117k 117k 117k 100k clipping factor {1, 0.95, 0.9} {1, 0.95, 0.9} {1, 0.95, 0.9} {1, 0.95, 0.9} {1, 0.95, 0.9}
@@ Appendix F Additional Experiments
Table 7: Additional experiments for Exp.1 with different bit-width configurations and different model sizes.
Bit-width Method ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSum ROUGE (Avg.)
@@
Table 8: Additional experiments for Exp.1 with different bit-width configurations and different model sizes.
Bit-width Method ROUGE-1 ROUGE-2 ROUGE-L ROUGE-LSum ROUGE (Avg.)
Qwen2.5-0.5B
BF16 Base 23.79 6.63 18.46 18.56 16.86
[…]
W4A4KV4 RTN 10.04 0.37 8.15 8.34 6.73
GPTQ 12.53 0.92 10.08 10.50 8.51
QuaRot 9.94 0.57 8.18 8.38 6.67
SpinQuant 12.16 1.22 10.69 10.72 8.70
STE 29.97 9.92 23.39 23.39 21.67
[…]
W4A8KV4 RTN 9.51 1.06 9.02 8.90 7.12
GPTQ 9.53 1.04 8.80 8.73 7.03
QuaRot 8.24 1.25 7.51 7.23 6.06
SpinQuant 9.10 1.11 8.31 8.12 6.66
STE 32.14 11.50 25.18 25.18 23.50
[…]
Qwen2.5-7B
BF16 Base 32.72 11.82 25.18 25.42 23.79
[…]
W4A4KV4 RTN 1.07 0.00 1.01 1.01 0.77
GPTQ 0.72 0.00 0.69 0.69 0.53
QuaRot 7.21 0.10 5.93 5.93 4.79
SpinQuant 6.87 0.29 5.97 6.12 4.81
STE 30.86 10.16 23.73 23.73 22.12
[…]
W4A8KV4 RTN 5.73 0.23 4.72 4.74 3.86
GPTQ 7.48 0.27 6.22 6.36 5.08
QuaRot 5.62 0.15 5.08 5.14 4.00
SpinQuant 0.64 0.30 5.64 5.81 4.54
STE 34.44 13.29 27.16 27.17 25.52
[…]
Table 9: Additional experiments for Exp.2 on different bit-width configurations and different model sizes.
Bit-width Method TruthfulQA MMLU-Pro BigBenchHard AGIEval GSM8K Math Avg.